OpenStudy (anonymous):

Please help: x/3x+8 - x+4/x+3 =1

OpenStudy (anonymous):

\[\frac{ x }{ 3x+8 }-\frac{ x+4 }{ x+3 }=1\]

OpenStudy (anonymous):

find the solutions

OpenStudy (anonymous):

i multiplied (3x+8)(x+4) -x(x+3)

OpenStudy (anonymous):

and multiplied 1 by the denominator

OpenStudy (anonymous):

i got 2x^2+23x+32=3x^2+17x+24

OpenStudy (anonymous):

hello???

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

almost, but no, here amplify by (x+3)(3x+8) x(x+3) - (x+4)(3x+8) = (x+3)(3x+8) \[x^2+3x-(3x^2+8x+12x+32)=3x^2+8x+9x+24\] by ordering \[5x^2+34x+56=0\] Now you solve the second grade equation (I'm gonna use this method, feel free to use factorization or wathever) \[\dfrac {-b \pm \sqrt{b^2-4ac}} {2a}=x1,x2\] for ax^2+bx+c=0 And you'll get -4 and - (14/5)

OpenStudy (anonymous):

thank u kind person u deserve a medal

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

I revised it, it gives -4 and -14/5, try using WolframAlpha to check it

OpenStudy (anonymous):

i got it thank u !!